A taxi leaves the station X for station y every 10 minutes .simultaneously , a taxi leaves the station Y also for station X for every 10 minutes. the taxi moves at the same constant speed and go from X to Y or vice versa in 2 hours. how many taxis coming from the other side will each taxi meet enroute from Y to X ?

Grade:12th pass

5 Answers

Ankit Jaiswal
7 years ago
Hi
I don’t know that we have to consider the taxi that our taxi will meet on the end points of its journey(that is station X amd Y). so here i am taking that we don’t have to consider those taxis
 
Now we can do this thing by 2 methods 
  1. mathematical 
  2. logical
I will like to go with logical one
See when the taxi is just about to leave there will be taxis comming from the opposite side 
our taxi will surely meet those taxis. so no. of taxis which will be i the path are 
[(total time each taxi takes to reach destination)/(total time after which each taxi leaves the station)] +1
in our case it is
\frac{120}{10}+1= 12+1=13
(+1 as we need to take both the last and the first taxi)
so we have total of 13 taxis on the path out of which we will subtract 1 as i am not considering the taxi meeting station Y(if u have to consider it don’t subtract 1)
now we need to calculate the no. of taxis which leave after our taxi leaves
thoso will be again
[(total time each taxi takes to reach destination)/(total time after which each taxi leaves the station)] +1
in our case it is
\frac{120}{10}+1= 12+1=13
(+1 as we need to take both the last and the first taxi)
now  we will subtract 2 we have counted the taxi twice which will leave station X when our taxi leaves station Y and also as i am not considering the taxi meeting station X(if u have to consider it don’t subtract 1)
so now we have a total of 12+11=
23 taxis
Ankit Jaiswal
7 years ago
 and 25 if u are considering the taxis that our taxi meet on the station X and Y
 
 
thanks
 
plz approve if u found this helpful
Narendra Sharma
5 years ago
There are 11 taxies leaving station x and go for Y.similarly there are 11 taxies reaching x . And one taxi already stand at starting on x station so 11+11+1=23 taxies
Abishek
5 years ago
See the first thing we need to know is that when 6 th taxi leaves the station x or y then 1st taxi fron x meets the 1st taxi from y ..so according to this we have 6 taxis from x meeting 6 taxis from y.so total of 12 taxis meet in the first hour.and the same in the second hour too...
12*2=24.... And then 1 taxi will not be able to meet any taxis at the end of 2 he so
24-1=23....Hence 23 is the answer......
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached answer to your problem below.
 
When the 6 th taxi leaves the station x or y then 1st taxi fron x meets the 1st taxi from y
So according to this we have 6 taxis from x meeting 6 taxis from y.so total of 12 taxis meet in the first hour.
and the same in the second hour too
Therefore, 12*2=24
But 1 taxi will not be able to meet any taxis at the end of 2 hr
so
24-1=23
Hence 23 is the answer
 
Hope it helps.
Thanks and regards,
Kushagra